Please help me find the inverse of functions with an example like this:
f(x)=8(x-13)

Answer by nerdybill(7384)

Please help me find the inverse of functions with an example like this:
f(x)=8(x-13)
replace f(x) with y:
y=8(x-13)
switch x and y and solve for y:
x=8(y-13)
x/8 = y-13
x/8 + 13 = y
.
Therefore, the inverse is:
x/8 + 13